what is x in the equation 3x squared + 2x =0
Answer by LagrangeSon678(5)

This problem can be solved by factoring:
more specifically by factoring out a common term.
In this case: 3x^2+2x=0, we can factor out an x, since it is common to both 3x^2 and 2x.

This results in: x(3x+2)=0.
Now we simply set each term to equal zero like so:
x=0
3x+2=0
Now solve for x,
Oviously, x=0, thats on solution
The second solution comes from 3x+2=0. First subtract 2 from both sides, giving us 3x=-2. Then divide both sides by 3, resulting in x=-2/3.
Our solution are :
x=0 and x=-2/3